A body starts with an initial velocity of 10m/s and moves along a straight line with a constant acceleration. When the velocity of the body is 50m/s the acceleration reverses in direction. Find the velocity of the body when it reaches the starting point?

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
As Given:
u=10,a,v=50
v=u+at
=>50=10+at
=>t=40/a

distance it travels till its acceleration gets reversed = S1
v2=u2+2as
=>2500=100+2aS1
=>S1=1200/a

distance it travels till its velocity becomes 0= S2
v2=u2+2as
=>2500=2xaxS2
=>S2=1250/a

Velocity at the end = ve
v2=u2+2as
=>ve2=2xax(1250/a +1200/a)
=>ve=70m/s
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)